What is the new volume when 100.0 mL of a gas at 780.0 mmHg is changed to 760.0 mmHg?
1
Expert's answer
2018-04-03T09:10:35-0400
p1V1=p2V2
V2 = P1V1/p2 = 100*780/760 = 102.63 (ml)